Railroad Crossing Accidents The data show the number of railroad crossing accidents for the 50 states of the United States for a specific year. Construct a histogram, frequency polygon, and ogive for the data. Comment on the skewness of the distribution. (The data in this exercise will be used for Exercise 15 in this section.)


1
Expert's answer
2021-06-07T01:31:49-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Class \ limits& Frequency \\ \hline
 1-43 & 24 \\
 \hdashline
 44-86 & 17 \\
 \hdashline
87-129 & 3 \\
 \hdashline
130-172 & 4 \\
 \hdashline
173-215 & 1 \\
 \hdashline
216-258 & 0 \\
 \hdashline
259-301 & 0 \\
 \hdashline
302-344 & 1 \\
 \hdashline
\end{array}"



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 Class & Class & Midpoint & Relative \\ 
limits & boundaries & & frequency \\
\hline
 1-43 & 0.5-43.5 & 22 & 0.48 \\
 \hdashline
 44-86 & 43.5-86.5 & 65 & 0.34 \\
 \hdashline
87-129 & 86.5-129.5 & 108 & 0.06 \\
 \hdashline
130-172 & 129.5-172.5 & 151 & 0.08 \\
 \hdashline
173-215 & 172.5-215.5 & 194 & 0.02 \\
 \hdashline
216-258 & 215.5-258.5 & 237 & 0 \\
 \hdashline
259-301 & 258.5-301.5 & 280 & 0 \\
 \hdashline
302-344 & 301.5-344.5 & 323 & 0.02 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Cumulative & Cumulative \\ 
 & frequency & relative frequency \\
\hline
 <0.5 & 0 & 0.00 \\
 \hdashline
 <43.5 & 24 & 0.48 \\
 \hdashline
<86.5 & 41 & 0.82 \\
 \hdashline
<129.5 & 44 & 0.88 \\
 \hdashline
<172.5 & 48 & 0.96 \\
 \hdashline
<215.5 & 49 & 0.98 \\
 \hdashline
<258.5 & 49 & 0.98 \\
 \hdashline
<301.5 & 49 & 0.98 \\
 \hdashline
<344.5 & 50 & 1.00 \\
 \hdashline
\end{array}"

Histogram for number of railroad crossing accidents




Frequency polygon for number of railroad crossing accidents




Ogive for number of railroad crossing accidents





Mean>mode. Right-skewed (positive) distribution.